OpenStudy (anonymous):

simplfy( 4x over y^3) ^-3

OpenStudy (anonymous):

solve for one variable at a time. start with y. (4x/y3)-3=0 +3 +3 (4x/y3)=3 now multiply both sides by 1/4x to cancel the 4x on the top. like this: (4x/3y) * 1/4x= 3/4x 3y=(3/4x) /3 /3 y=(3/4x)/3 now take this and plug it back into the original problem for the y value to get your x. like this: [(4x) / {(3/4x)/3} * 3] -3. now solve that to get your x. then take your x value from that and plug it in to the original again to get your actual y value.

OpenStudy (mathstudent55):